OpenStudy (anonymous):

if f(x) = x/(x^2-16) then what is the value of f(-10)?

jimthompson5910 (jim_thompson5910):

f(x) = x/(x^2-16) f(-10) = -10/((-10)^2-16) f(-10) = -10/(100-16) f(-10) = -10/(84) f(-10) = -5/42